Find all real or imaginary solutions to each equation.
x^-2 - 9x^-1 + 18 = 0

Answer by stanbon(75887)

x^-2 - 9x^-1 + 18 =0
(x^(-1)-6)(x^(-1)-3)=0
x^-1=6 or x^-1=3
x=1/6 or x=1/3
Cheers,
Stan H.